OpenStudy (mef4.0):

find the derivative of: f(x) = (sec x)(x^2 +1)

OpenStudy (anonymous):

All you need to do is use the rules of integration. So it is (secx)'(x^2 + 1) + (x^2 +1)(secx)' See if you can finish it off

OpenStudy (mef4.0):

Sorry. I don't understand your notation - are you telling me to use the product rule?

OpenStudy (anonymous):

Sorry Yes you want to use the product rule here sec'(x) means the derivative of sec(x) (x^2 +1)' means the derivative of x^2 +1 It may help to think of the problem like this h(x) = sec(x) g(x) = x^2 + 1 f(x) = h(x)g(x)